From mar@roe.ac.uk Mon Nov 17 11:58:25 2014 Date: Mon, 17 Nov 2014 10:24:29 +0000 From: Mike Read <mar@roe.ac.uk> To: VDFS Science Archive Team -- Clive Davenhall <acd@roe.ac.uk>, Dave Morris <dmr@roe.ac.uk>, Eckhard Sutorius <etws@roe.ac.uk>, Mark Holliman <msh@roe.ac.uk>, Nicholas Cross <njc@roe.ac.uk>, Nigel Hambly <nch@roe.ac.uk>, Rob Blake <rpb@roe.ac.uk>, Ross Collins <rsc@roe.ac.uk>, Stelios Voutsinas <stv@roe.ac.uk> Cc: CCs for VDFS -- Alastair Edge <Alastair.Edge@durham.ac.uk>, Andrew Lawrence <al@roe.ac.uk>, Bob Mann <rgm@roe.ac.uk>, Dante Minniti <dante@astro.puc.cl>, James Dunlop <jsd@roe.ac.uk>, Jim Emerson <j.p.emerson@qmul.ac.uk>, Keith Noddle <keith@keithnoddle.org>, Lorenzo Rimoldini <Lorenzo.Rimoldini@unige.ch>, Maria-Rosa Cioni <mcioni@aip.de>, matt.jarvis@astro.ox.ac.uk, Mike Irwin <mike@ast.cam.ac.uk>, Nigel Metcalfe <nigel.metcalfe@durham.ac.uk>, Norman Gray <norman@astro.gla.ac.uk>, Phil Lucas <pwl@star.herts.ac.uk>, Richard McMahon at IOA <rgm@ast.cam.ac.uk>, Simon Dye <Simon.Dye@nottingham.ac.uk>, Stephen Warren <s.j.warren@ic.ac.uk>, Tom Kerr <tkerr@jach.hawaii.edu>, Tom Shanks <tom.shanks@durham.ac.uk> Subject: WFAU VDFS Science Archives progress meeting minutes, 14 November 2014 WFAU VDFS Science Archives progress meeting minutes, 14 November 2014 Present: MAR, STV, ETWS, RSC, ACD, NJC, MSH Apologies: DM, NCH, KTN, AL, RGM, RPB DoNM: 10am Friday 28 November 2014 in the Villas meeting room, ACD in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: RSC to contact Iraklis Klampanos to discuss performance of alternative RDBMSs. Discharged (though no reply and became unnecessary once we discovered an index was missing on a VVV table). Actions carried forward from 17/10/2014 meeting: ---------------------------------------------- ACTION: RPB (was MAR) to have a look at the status of non-survey releases for registered projects. MAR had compiled a list of non-survey projects that could be released and sent it to RPB. RPB had yet to make these data available. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. However, the action continues as RGM had not yet established which early releases could safely be scrapped. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this action still has low priority. ACTION: All to finish off VVV brainstorming tasks by end of January at the latest. Continues. ACTION: MSH to re-configure RAID on /sahure. Continues, data has been copied off. ACTION: ACD speak to HME about the possibility of IT support making nightly backups of the SVN repository. Continues. ACTION: RPB write a chron job to make daily check-outs of all the WFAU Projects in the SVN repository in order to provide a backup snapshot of the current state. Continues. Specific points and new actions: ------------------------------- Project management: Nothing for the minutes WFCAM, VISTA and VST updates: The UKIRT handover ceremony had taken place. Review of Deadlines: ACD noted that he would be re-working the deadline page and the result would form the basis of future reviews. ACTION: ACD to re-vamp the deadlines page. Networking: Working. WSA/VSA/OSA/GES Operations: ETWS reported that OSA ingests were still running. Some list driven catalogues have missing header information eg APERCOR values, resulting in failed ingests. VVV files identified as v1.3.1 had been transferred and would need to be ingested for DR3. The available WFCAM satellite data had been CU3'ed. ACD noted that additional GES iDR3 files had been copied over. Hardware and Systems: All running. Software: A new version of mxODBC was available, the consensus was to give it a try as it might provide performance benefits. ACTION: ETWS/NCH purchase license for new mxODBC. Survey Data Release: UKIDSS DR10 had been made public. The inclusion of GPS in proprietary DR10 was still pending Phil Lucas's approval. There had been VIDEO consortium and ESO releases. Non-survey Data Release: A database release had been made for U13A04. Astrogrid deployment, VO & Data Analysis services: STV reported that he and DM had continued work on extending ADQL and distributed queries using OGSA-DAI. Miscellaneous: ACD asked if all those contacted about checking their travel claims could please do so as soon as possible. Upcoming meetings: Tools for Astronomical Big Data - Tucson, Arizona, March 9-11, 2015 VMC meeting - Naples July 2015 VVV meeting - Albano May? 2015 -- Scanned by iCritical.